Assume out of 500 people 280 are rice eater and rest are wheat eater. Can
we say the proportion of rice and wheat eater are same at 95%?
1
Expert's answer
2020-06-11T19:27:02-0400

For rice, we have that the sample size is "n_1=500," the number of favorable cases is "x_1=280," so then the sample proportion is "\hat{p}_1=\dfrac{280}{500}=0.56"

For wheat, we have that the sample size is "n_2=500," the number of favorable cases is "x_2=220," then the sample proportion is "\hat{p}_2=\dfrac{220}{500}=0.44"

The value of the pooled proportion is computed as


"\bar{p}={x_1+x_2\over n_1+n_ 2}=0.5"

Also, the given significance level is "\alpha=0.05"

The following null and alternative hypotheses need to be tested:

"H_0:p_1=p_2"

"H_1:p_1\not=p_2"

This corresponds to a two-tailed test, for which a z-test for two population proportions needs to be conducted.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "z_c=1.96"

The rejection region for this two-tailed test is "R=\{z:|z|>1,96\}"

The z-statistic is computed as follows:


"z={\hat{p}_1-\hat{p}_2 \over \sqrt{\bar{p}(1-\bar{p})(1/n_1+1/n_2)}}="

"={0.56-0.44 \over \sqrt{0.5(1-0.5)(1/500+1/500)}}=3.795"

Since it is observed that "|z|=3.795>1.96=z_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that population proportion "p_1" is different than "p_2"​, at the 0.05 significance level.

Using the P-value approach: The p-value is "p=0.000148," and since "p=0.000148<0.05," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that population proportion "p_1" is different than "p_2", at the 0.05 significance level.